amelyajankens:

A supervisor finds the mean number of miles that the employees in a department live from work. He finds mc020-1.jpg and mc020-2.jpg. Which mileage is within a z-score of 1.5?

Tranquility:

Welcome to QuestionCove! Please provide the images by using the "attach file" button

Tranquility:

@amelyajankens

Evo:

this is question: A supervisor finds the mean number of miles that the employees in a department live from work. He finds x=29 and s=3.6. Which mileage is within a z-score of 1.5?

Ferredoxin4:

z-score is basically another term for the specific number of standard deviations. I suppose we can assume normality here, even though it isn't given, so I'd make a rough sketch. Otherwise, just use the basic formula: \[Z_{score} = \frac{ X-μ }{ σ }\] It's given that x̄=29 (statistic for a mean) and s=3.6 (statistic for a standard deviation. Z-score must be LESS than 1.5 \[|1.5>\frac{ x-29 }{ 3.6 }|\] Simple algebra here now. I just used normalcdf and got 23.6<x<34.4 I'm not sure what your options are, so just match them.

Ferredoxin4:

 Just a diagram so you understand it better.